Estradiol valerate is an estrogen drug used to treat menopausal vasomotor symptoms. A typical dose, administered by deep IM injection, consists of 20 mg once every four weeks. If a vial of estradiol valerate contains 5 ml with a concentration of 10 mg/ml, how many vials will be needed for 40 weeks of therapy?

Follitropin alfa is a fertility drug that stimulates follicle maturation and induces ovulation. A typical dose consists of subcutaneous injections of 75 units/day for 14 days. If 100 units equals 1 milliliter, how many milliliters will be needed for a 2-week treatment?

For the treatment of amenorrhea, norethindrone is administered orally in doses of up to 10 mg/day for 10 consecutive days during the second half of the menstrual cycle. If 5-mg tablets of norethindrone are administered throughout the regimen, how many tablets will be needed?

Oxymethalone is an anabolic steroid used to treat anemia. An effective dose, which is administered over a 3- to 6-month period, is typically 2 mg/day for every kilogram of body mass. For a patient whose weight is 192.5 pounds, how many 50-mg tablets are needed for a daily dose of oxymethalone?
Answer #1

1.Estradiol valerate dose = 20 mg once every four weeks.

If a vial of estradiol valerate contains 5 ml with a concentration of 10 mg/ml,

Hence for 20 mg dose required ml= 2ml, for 4 weeks

Then for 40 weeks required ml=20 ml

The no. of vials needed for 40 weeks of therapy= 20/5=4 vial of estradiol valerate contains 5 ml will be needed.

2.

Follitropin alfa typical dose consists of subcutaneous injections of 75 units/day for 14 days.

then for 2 weeks required units= 75X14=1050 units

If 100 units = 1 ml

Required ml for a 2-week treatment?

100 units=1ml

1 unit=1/100ml

1050 units=1050/100=10.5 ml will be required for 2 weeks treatment

3.

For the treatment of amenorrhea, norethindrone is administered orally in doses =10 mg/day for 10 consecutive days.

Total amount in dose taken for 10 days =100 mg

If 5-mg tablets of norethindrone are administered throughout the regimen

How many tablets will be needed?

=100/5=20 tablets

4. Oxymethalone is an anabolic steroid used to treat anemia. An effective dose, which is administered over a 3- to 6-month period= 2 mg/day for every kilogram of body mass.

For a patient whose weight is 192.5 pounds=87.5 kg

Hence daily dose for the 87.5 kg weight patient=175mg per day

How many 50-mg tablets are needed for a daily dose of oxymethalone?

50mg in 1 tablet

hence 175mg will be supplied by =175/50=3.5 tablets.